OpenStudy (anonymous):

What is 240 16% and 20% of? and how do you figure it out?

OpenStudy (anonymous):

38.4

OpenStudy (iwillrektyou):

Hi, do you know how to find a portion of a number? For example, 25% of is 200 is 50.

OpenStudy (anonymous):

48

OpenStudy (iwillrektyou):

To do this, you just do the inverse operation. I will try to guide you through this so you can understand how to do it.

OpenStudy (iwillrektyou):

To start off with the basics, to find 25% of 200, you just multiply: 200 * 0.25 = 50.

OpenStudy (anonymous):

I know how to do that...

OpenStudy (iwillrektyou):

Ok. So in order to find this, you have to divide.

OpenStudy (iwillrektyou):

Do you have a question on that?

OpenStudy (texaschic101):

do you understand ?

OpenStudy (anonymous):

no i don't

OpenStudy (iwillrektyou):

Ok, so in order to solve, you have to do the same thing of finding the portion of 200, except you have to divide 0.16 and 0.2

OpenStudy (texaschic101):

Read your problem, then write it like you read it. just remember, in " math language ", ---- " of " means multiply and " is " means equals. for example : 25% of what is 200 first change % to decimal.. 0.25 * x = 200 -- divide both sides by 0.25 x = 200/0.25 x = 800 or 16% of 200 is what number 0.16 * 200 = x 32 = x or you can just use the percent formula... is/of = %/100 and if you get this problem... 25% of what is 200 % = 25 of = x is = 200 now sub into the formula... 200/x = 25/100 this is a proportion, so you cross multiply 25x = 20000 x = 20000/25 x = 800